# Usable XAI: 10 Strategies Towards Exploiting Explainability in the LLM Era

### v1

@@ 4.2.2 Results and Analysis
Table 4: Effectiveness of EK-FAC approximated influence function on the established SciFact-Inf dataset. Time (Pre.) stands for the time for precomputing the $\mathbf{Q}_{A}$, $\mathbf{Q}_{S}$, and $\bm{\Lambda}$. Time (Inf.) stands for the time for calculating the influence of 100 training samples per test sample. Experiments for GPT2-1.5B, LLaMA2-7B, Mistral-7B are done on 4 A100 GPUs, whereas experiments for LLaMA2-13B are done with 8 A100 GPUs.
@@
The experimental results are summarized in Table 4. From Table 4 we can find that, the EK-FAC approximated influence function achieves a good accuracy in finding the training sample that has the greatest influence on the generation of a test sample, even if only the influences mediated by a small part of dense layers are considered. In addition, we find that the main computational bottleneck in calculating the EK-FAC-based influence is to estimate the covariances $\mathbf{Q}^{(l)}_{A}$, $\mathbf{Q}^{(l)}_{S}$ and variance $\bm{\Lambda}^{(l)}$, which can take hours when 500 training samples are used for the estimation. However, after the estimation, it is relatively cheap to calculate the influence of 500 training samples for each test sample, which takes around 3 minutes for a 13B LLaMA2 model on 8 A100 GPUs. This demonstrates the scalability of EK-FAC-based influence by assuming the independence of different dense layers and using EK-FAC to simplify the computation.
4.3 Challenges
@@
Overall, explaining the generation of LLMs by tracing back to the training samples is still an emerging area. Open questions need to be addressed to further advance the field. In this section, we identify three main challenges as follows, which can serve as directions for future explorations.
4.3.1 Strong Assumptions for Scalability
@@
The unprecedented number of parameters in modern LLMs causes severe scalability issues for sample-based explanation strategies. This is especially evident for the gradient-based methods, as the HVP in Eq. (1) induces both high computational and space complexity. To address the bottleneck, strong assumptions are usually required to make it feasible for large models. For example, TracIn (Pruthi et al., 2020) simplifies the second-order term in Eq. (1) via first-order approximation. Schioppa et al. (2022) assume the Hessian to be low rank. Grosse et al. (2023) that assume that the weights from different layers of the LLMs are independent, as well as the tokens in different steps, such that EK-FAC can be appropriately applied to approximate the influence function. From the above analysis, we can find that while the method from Grosse et al. (2023) has the best scalability, it also has the strongest assumption, which may fail to hold in practice. While highly efficient to compute, embedding-based methods make the implicit assumption that semantics similarly implies explainability, which we have demonstrated may not always be the case. Therefore, how to improve the scalability with weak assumptions needs to be investigated in the future.
4.3.2 Explainability v.s. Understandability
Despite the advantage of influence/embedding similarity to provide a qualitative measurement of a specific training sample as the explanation for LLM generation, the understandability of the identified sample can still be weak, where the connection between the selected training samples and the generation may not be understandable to human beings. Specifically, Grosse et al. (2023) cautions that the sign of influence score of the training tokens may be difficult for humans to connect to the positive or negative influence on the generation results. This severely jeopardizes the usability of the identified training samples. In addition, Grosse et al. (2023) also found that, since LLMs are usually not trained to the minimum to avoid overfitting (and due to overparameterization, the number of local minimums may be large), the connection between influence defined in Eq. (1) with the counterfactual loss of removing the sample $z_{i}$ at $z$ is also weak. For the embedding-based methods, since most LLM models are black box transformer models, the similarity of embeddings can also be hard to interpret by human beings; therefore, it is imperative to improve the interpretability of the identified training samples, such that tracing back becomes more meaningful.
4.3.3 LLM-Oriented Sample-based Explanations
@@
Finally, we observed that both gradient-based and embedding-based methods are loosely connected to the LLM, as well as the backbone transformer networks. For example, algorithms like TracIn (Pruthi et al., 2020) are designed to scale up influence functions to large models, which are not specific for LLMs. Similarly, the embedding-based method proposed in Akyurek et al. (2022) is applicable to most machine learning models with latent representations. Grosse et al. (2023) considers the specialty of LLMs by utilizing the knowledge neuron assumption of the backbone transformers (Wang et al., 2023d) to simplify the influence function, where the weights considered are constrained to the MLP layers, which may not fully utilize the property of transformers. Therefore, how to further utilize the property of the LLM and the backbone transformer to design LLM-tailored sample-based influence/similarity (either to reduce the computational/space overhead or to improve the explanation quality) is highly promising for future work.
5 Explainability for Trustworthy LLMs and Human Alignment
@@
In previous sections, we explore the use of explanation techniques for assessing and improving the performance of LLMs. In this section, we shift the focus towards examining LLM trustworthiness. As LLMs are increasingly integrated into various applications of daily life, including high-stakes areas like healthcare, finance, and legal advice, it is crucial that their responses not only are accurate but also align with human ethical standards and safety protocols (Liu et al., 2023b; Li et al., 2023f). Thus, the need arises to extend the scope of explanations from merely evaluating the accuracy of LLMs to scrutinizing their trustworthiness. Herein, we delve into how explanation techniques, discussed in the previous sections, can be instrumental in assessing LLMs across key aspects of trustworthiness like security, privacy, fairness, toxicity, and honesty. It is worth noting that while explainability itself is an aspect of trustworthiness, it holds the promise of serving as a foundational tool for addressing other trustworthiness concerns.
5.1 Security
@@
LLMs are known to be vulnerable to attacks and exploitation, such as spreading misinformation, launching phishing attacks, and poisoning training data (Derner et al., 2023). For enhanced safety, LLMs are designed to reject certain types of prompts that may result in the generation of harmful content, e.g., by precluding prompts likely to elicit unsafe outputs from the training phase. However, jailbreak techniques can circumvent these restriction measures and manipulate LLMs into producing malicious contents. Malevolent users (i.e., attackers) can craft special prompts that compel or induce LLMs to prioritize instruction following over rejections (Liu et al., 2023c; Li et al., 2023a). For example, through Prefix Injection, attackers can use out-of-distribution prompt prefixes that are less likely to be rejected (walkerspider, 2022; Wei et al., 2023). Another approach, called Refuse Suppression, involves directing or persuading models to ignore established safety protocols (Wei et al., 2023; Zeng et al., 2024), where the instruction following ability is then employed to perform the attack.
@@
Existing methods mainly rely on prompt engineering for attacking LLMs, but they usually have low attack success rates and significant time costs (Li et al., 2024c). Thus, by understanding and engineering latent representations of LLMs, explanation methods provide a viable way to design advanced attacks and discover the potential vulnerabilities of LLMs (Liu et al., 2021). For example, a recent work extracts “safety patterns” via explaining the latent space of LLMs using representation engineering. Specifically, these patterns can be captured from the activation differences between malicious queries and benign queries. The salient portion of difference vectors’ dimensions is localized and utilized to generate features of safety patterns. The safety patterns reflect the internal protection mechanisms within LLMs. Circumventing these patterns leads to novel attacks, which helps exploring potential vulnerabilities of LLMs (Li et al., 2024c). Besides, a deeper understanding of fine-tuning can shed light on the reliability of existing safety measures. In particular, Jain et al. (2023) use networking pruning, attention map activation, and probing classifiers to track the changes of model capabilities from pre-training to fine-tuning. These tools are helpful in finding salient weights that identify crucial neurons to restore pre-training capabilities. These neurons have demonstrated that the capabilities gained during fine-tuning can be removed easily through fine-tuning on other unrelated tasks. This finding casts doubt on the robustness of current safety alignments in LLMs.
5.2 Privacy
@@
Recent studies have revealed that LLMs such as ChatGPT can leak extensive amounts of training data through a method known as divergence attacks. These attacks utilize specially crafted prompts to lead the model away from its standard chatbot-style generation (Nasr et al., 2023). The risk of private data exposure through such means poses a serious challenge to the development of ethically responsible models. This issue is compounded by strategies similar to jailbreak attacks, where misalignment is exploited to induce LLMs into operating in an unconventional “developer mode” via out-of-distribution prompts (Li et al., 2023a). Traditional data pre-processing techniques, such as data sanitization (Lison et al., 2021), are impractical as defenses for LLMs due to the enormous scale of training data.
@@
Enhancing LLM privacy involves two strategic approaches: (1) preventing the models from memorizing sensitive data, and (2) establishing safeguards against the release of sensitive information during content generation. The latter can employ techniques used in jailbreak defenses, treating prompts that solicit private information as potentially malicious. The former approach requires identifying whether LLMs possess specific knowledge, which is traditionally tested by crafting question-answering (QA) tasks to determine if LLMs can provide answers. However, this approach faces limitations due to LLMs’ sensitivity to the phrasing of QA prompts, while the optimal prompt is usually unknown. To tackle the challenge, explanatory techniques can serve as a tool to confirm whether LLMs have internalized certain knowledge. For instance, via explaining the relation between factual knowledge and neuron activations (Meng et al., 2022a; Dai et al., 2022; Hase et al., 2024), we may investigate whether and where a piece of factual knowledge is stored within transformers. In addition, Yin et al. (2024a) recently proposes the concept of “knowledge boundary” and develops a gradient-based method to explore whether LLMs master certain knowledge independent of the input prompt.
5.3 Fairness
Despite LLMs’ powerful generation capabilities, their widespread applications also bring concerns about exacerbating bias issues in society, as LLMs are able to learn social biases within human-generated corpus (Gallegos et al., 2023). For example, in a gender bias case, “[He] is a doctor” is much more likely than “[She] is a doctor” because of the gender stereotype. In this subsection, we focus on fairness issues that refer to biases related to race, gender, and age within human communities (Li et al., 2023g). There is a rich body of literature quantifying fairness issues within LLMs through various tests (Adebayo et al., 2023). Interpretation complements these methods by providing a unique perspective for bias mitigation, focusing on unraveling the mechanisms through which biases are embedded into LLMs. A research direction within this domain is the examination of biased attention heads. For instance, Ma et al. (2023) detect stereotype encodings through probing attention heads and evaluating their attributions via Shapley values (Lundberg & Lee, 2017). The results unveil that approximately 15% to 30% of attention heads across six transformer-based models are linked to stereotypes. These attention heads tend to specialize in maintaining various stereotypes, offering a potential pathway for developing effective de-biasing techniques. Moreover, scoring head biases with gradient-based metrics offers another way for identifying biased heads (Yang et al., 2023a). Furthermore, recent work has placed LLM representations under scrutiny (Zou et al., 2023a). Typically, specific templates relevant to a given concept or function are designed beforehand. Then, representations closely aligned with the concepts or functions are examined using principal component analysis (PCA). From this analysis, a vector is derived from the first principal component to predict a certain bias.
@@ 5.3 Fairness
To achieve fair model predictions, a diverse range of mitigation techniques have been proposed to debias models. One stream of work proposes to debias LLMs at the embedding level, refining embeddings through various approaches. For example, a recent work attempts to alter biased embeddings with minimal alterations to make them orthogonal to neutral embeddings (Rakshit et al., 2024). Additionally, some studies concentrate on removing biases at the level of attention heads. Ma et al. (2023) address this by pruning attention heads that significantly contribute to certain biases. Similarly, another study adapts movement pruning (Sanh et al., 2020) to dynamically choose an optimal subset of low-bias attention heads (Joniak & Aizawa, 2022). Beyond modifying embeddings and pruning attention heads, another strategy involves targeting a specific group of neurons known to propagate biases. It unlearns the biases by retraining weight vectors for these neurons (Yu et al., 2023). Besides, bias mitigation can also be approached from a data-centric perspective using a few training samples (Thakur et al., 2023). This work uses a pre-trained model to find the most biased training examples, and then modifies these examples to fine-tune the model.
5.4 Toxicity
@@
Toxicity is another form of harmful content that LLMs may produce. This issue arises because LLMs are often trained on extensive online text corpora that have not been thoroughly filtered, containing elements of toxicity that can hardly be fully eliminated. Toxicity can be identified by interpreting LLM components like the feed-forward layers and attention heads. For instance, recent work reveals how toxicity is represented within LLMs by identifying multiple vectors promoting toxicity within the MLP layers and further recognizes relevant dimensions through singular value decomposition (Lee et al., 2024). Furthermore, the exploration of geometric structures in per-layer representations offers another way to detect toxicity. Balestriero et al. (2023) applied spline formulations within MLPs to extract seven characteristic input features, demonstrating their utility in describing the domain of prompts and classifying toxic remarks.
@@
The insights into how toxicity manifests within LLMs also shed light on mitigation strategies. Motivated by the finding that toxicity can be reduced by manipulating relevant vectors, Lee et al. (2024) develops a method called direct preference optimization (DPO), where paired toxic and non-toxic samples are utilized to fine-tune models so that non-toxic content is promoted. By examining the changes in the parameter matrices during the fine-tuning process, it substantiates that even minor adjustments to these critical vectors can reduce toxicity. Built on the observation that LLMs’ representations are updated by outputs from attention layers (Elhage et al., 2021), another work attempts to reduce toxicity by identifying the “toxicity direction” and then adjusting representations in the opposite direction (Leong et al., 2023).
5.5 Truthfulness

### v2

@@ 4.2.2. Results and Analysis
Table 2. Effectiveness of EK-FAC approximated influence function on the SciFact-Inf dataset. Time (Pre.) stands for the time for precomputing the $\mathbf{Q}_{A}$, $\mathbf{Q}_{S}$, and $\boldsymbol{\Lambda}$. Time (Inf.) stands for time for calculating the influence of 100 training samples per test sample.
@@
The results are summarized in Table 2. We observe that the EK-FAC approximated influence function achieves good accuracy in finding the training sample with greatest influence on the generation of LLMs. In addition, we find that the main computational bottleneck in calculating the EK-FAC-based influence is to estimate the covariances $\mathbf{Q}^{(l)}_{A}$, $\mathbf{Q}^{(l)}_{S}$ and variance $\boldsymbol{\Lambda}^{(l)}$, which takes hours when 500 training samples are used for the estimation. However, after that, it is relatively cheap to calculate the influence training samples for each test sample.
4.3. Challenges
@@
Overall, explaining the generation of LLMs by tracing back to the training samples is still an emerging area. Open questions need to be addressed to further advance the field.
4.3.1. Strong Assumptions for Scalability
@@
The unprecedented number of parameters in modern LLMs causes severe scalability issues for sample-based explanation strategies. This is especially evident for the gradient-based methods, as the HVP in Eq. (8) induces both high computational and space complexity. To address the bottleneck, strong assumptions are usually required to make it feasible for large models. For example, TracIn (Pruthi et al., 2020) simplifies the second-order term in Eq. (8) via first-order approximation. (Schioppa et al., 2022) assume the Hessian to be low rank. (Grosse et al., 2023) that assume that the weights from different layers of the LLMs are independent, as well as the tokens in different steps, such that EK-FAC can be appropriately applied to approximate the influence function. From the above analysis, we can find that while the method from (Grosse et al., 2023) has the best scalability, it also has the strongest assumption, which may fail to hold in practice. Therefore, how to improve the scalability with weak assumptions needs to be investigated in the future.
4.3.2. LLM-Oriented Sample-based Explanations
@@
Finally, we observed that both gradient-based and embedding-based methods are loosely connected to the LLM, as well as the backbone transformer networks. For example, algorithms like TracIn (Pruthi et al., 2020) are designed to scale up influence functions to large models, which are not specific for LLMs. Similarly, the embedding-based method proposed in (Akyurek et al., 2022) is applicable to most machine learning models with latent representations. (Grosse et al., 2023) considers the specialty of LLMs by utilizing the knowledge neuron assumption of the backbone transformers (Wang et al., 2023d) to simplify the influence function, where the weights considered are constrained to the MLP layers, which may not fully utilize the property of transformers. Therefore, how to further utilize the property of the LLM and the backbone transformer to design LLM-tailored sample-based influence/similarity (either to reduce the computational/space overhead or to improve the explanation quality) is highly promising for future work.
5. Explainability for Trustworthy LLMs and Human Alignment
@@
In previous sections, we explore the use of explanation techniques for assessing and improving the performance of LLMs. In this section, we shift the focus towards examining LLM trustworthiness. As LLMs are increasingly integrated into high-stakes areas like healthcare, finance, and legal advice, it is crucial that their responses not only are accurate but also align with human ethical standards and safety protocols (Liu et al., 2023c; Li et al., 2023a; Hu et al., 2023). Herein, we examine how explanation techniques, discussed in the previous sections, can be instrumental in assessing LLMs across key aspects of trustworthiness like security, privacy, fairness, toxicity, and honesty. It is worth noting that while explainability itself is an aspect of trustworthiness, it holds the promise of serving as a foundational tool for addressing other trustworthiness concerns.
5.1. Security
@@
LLMs are known to be vulnerable to attacks and exploitation, such as spreading misinformation, and poisoning training data (Derner et al., 2023). For enhanced safety, LLMs are designed to reject prompts that may result in harmful contents. However, jailbreak techniques can circumvent these restriction measures and manipulate LLMs into producing malicious contents. Malevolent users can craft special prompts that compel LLMs to prioritize instruction following over rejections (Liu et al., 2023a; Li et al., 2023c). For example, through Prefix Injection, attackers can use out-of-distribution prompt prefixes that are less likely to be rejected (walkerspider, 2022; Wei et al., 2023). Another approach, called Refuse Suppression, involves directing or persuading models to ignore established safety protocols (Wei et al., 2023; Zeng et al., 2024a), where the instruction following ability is then employed to perform the attack.
{forest}
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, rectangle, draw, font=, rounded corners, align=center, minimum width=2em, edge+=darkgray, line width=1pt, s sep=4pt, inner xsep=3pt, inner ysep=2pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, text centered, , [Trustworthy LLMs & Human Alignment,ver,fill=gray!20,text width=10em,text centered [Security,ver,fill=orange!40,text centered,text width=2.5em [Adversarial Attack,fill=orange!10,text centered,text width=6.5em [(Liu et al., 2023a), (Li et al., 2023c), (Wei et al., 2023), (Zeng et al., 2024a), (walkerspider, 2022), (Liu et al., 2021), (Jain et al., 2023)., leaf1, align=left] ] [Adversarial Defense,fill=orange!10,text centered,text width=6.5em [(Wu et al., 2025c), (Yu et al., 2024b), (Zheng et al., [n. d.]), (Yousefpour et al., 2025)., leaf1, align=left] ] ] [Privacy,ver,fill=orange!40,text width=2.5em,text centered [Memorization Attack,fill=orange!10,text centered,text width=6.5em [(Nasr et al., 2023), (Li et al., 2023c), (Fu et al., 2024), (Bai et al., 2024)., leaf1, align=left] ] [Knowledge-based Defense,fill=orange!10,text centered,text width=6.5em [(Meng et al., 2022a), (Dai et al., 2022), (Hase et al., 2024), (Yin et al., 2024b), (Suri et al., 2025)., leaf1, align=left] ] ] [Fairness,ver,fill=orange!40,text width=2.5em,text centered [Bias Detection,fill=orange!10,text centered,text width=6.5em [(Adebayo et al., 2023), (Ma et al., 2023), (Zou et al., 2023a), (Qian et al., 2024)., leaf1, align=left] ] [Bias Mitigation,fill=orange!10,text centered,text width=6.5em [(Rakshit et al., 2024), (Ma et al., 2023), (Yu et al., 2023), (Thakur et al., 2023)., leaf1, align=left] ] ] [Toxicity,ver,fill=orange!40,text width=2.5em,text centered [Toxic Content Detection,fill=orange!10,text centered,text width=6.5em [(Balestriero et al., 2023), (Wang et al., 2024d), (Lee et al., 2024), (Li et al., 2024d)., leaf1, align=left] ] [Toxic Content Mitigation,fill=orange!10,text centered,text width=6.5em [(Lee et al., 2024), (Liu et al., 2024b), (Leong et al., 2023)., leaf1, align=left] ] ] [Truthfulness,ver,fill=orange!40,text width=2.5em,text centered [Hallucination Detection,fill=orange!10,text centered,text width=6.5em [(Azaria and Mitchell, 2023), (Campbell et al., 2023), (Duan et al., 2024), (Li et al., 2024b), (Marks and Tegmark, 2023), (Xu et al., 2023a), (Heo et al., 2024), (Park et al., 2025)., leaf1, align=left] ] [Hallucination Mitigation,fill=orange!10,text centered,text width=6.5em [(Li et al., 2024b), (Chen et al., [n. d.]), (Liu et al., 2024b), (Yu et al., 2024a), (Huang et al., 2024), (Zhang et al., 2024c), (He et al., 2024)., leaf1, align=left] ] ] ]
Figure 4. Taxonomy of explanation techniques for trustworthiness aspects of LLMs.
@@
By engineering latent representations of LLMs, explanation methods provide a viable way to discover the potential vulnerabilities of LLMs by simulating advanced attacks (Liu et al., 2021). For example, a recent work extracts “safety patterns” via explaining the latent space. These patterns are captured from the activation differences between malicious and benign queries. They reflect the internal protection mechanisms within LLMs. Circumventing these patterns leads to novel attacks, which helps exploring potential vulnerabilities of LLMs (Li et al., 2024d). Besides, a deeper understanding of fine-tuning can shed light on the reliability of existing safety measures. In particular, Jain et al. (2023) apply network pruning, attention maps, and probing classifiers to track the changes of model capabilities from pre-training to fine-tuning. These tools are helpful in demonstrating that the capabilities gained during fine-tuning can be easily removed through fine-tuning on other unrelated tasks. This finding casts doubt on the robustness of current safety alignments in LLMs. Finally, by probing LLM representations, Zheng et al. ([n. d.]) investigates the impact of safety prompts that motivates a novel safety prompt optimization method to safeguard LLMs.
5.2. Privacy
@@
Recent studies have revealed that LLMs such as ChatGPT can leak extensive amounts of training data through divergence attacks (Nasr et al., 2023). These attacks utilize specially crafted prompts to lead the model away from its standard chatbot-style generation. The risk of private data exposure through such means poses a serious challenge to the development of ethically responsible models. This issue is compounded by strategies similar to jailbreak attacks, where misalignment is exploited to induce LLMs into operating in an unconventional “developer mode” via out-of-distribution prompts (Li et al., 2023c).
@@
Enhancing LLM privacy involves two strategic approaches: (1) preventing LLMs from memorizing sensitive data, and (2) establishing safeguards against the release of sensitive information. The second strategy involves employing techniques from jailbreak defenses, treating prompts that solicit private information as potentially malicious. The first approach requires identifying specific knowledge within models, which is traditionally achieved with prompt engineering. However, prompt engineering faces limitations due to LLMs’ sensitivity to the phrasing of prompts. Explanatory techniques can serve as an auxiliary tool to confirm whether LLMs have internalized certain knowledge. For instance, factual knowledge is localized via explaining the relation between factual knowledge and neuron activations (Meng et al., 2022a; Dai et al., 2022; Hase et al., 2024). In addition, (Yin et al., 2024b) recently proposes the concept of “knowledge boundary” and develops a gradient-based method to explore whether LLMs master certain knowledge independent of the input prompt.
5.3. Fairness
The widespread applications of LLMs also bring concerns about exacerbating bias issues in society (Gallegos et al., 2023). For example, in a gender stereotype case, “[He] is a doctor” is much more likely than “[She] is a doctor”. In this subsection, we focus on fairness issues related to race, gender, and age (Li et al., 2023b). Prompt engineering is a major method to quantify fairness issues within LLMs (Adebayo et al., 2023). Interpretation complements this method by unraveling the mechanisms through which biases are embedded into LLMs. One direction is examining biased attention heads. Ma et al. (2023) unveil that approximately 15% to 30% of attention heads linking to specific stereotypes through probing attention heads. Furthermore, recent work has placed LLM representations under scrutiny (Zou et al., 2023a). Typically, representations of crafted templates showing specific biases are examined to derive a vector that identifies a certain bias.
@@ 5.3. Fairness
To achieve fair model outputs, a diverse range of mitigation techniques have been proposed. One stream of work proposes to debias LLMs at the embedding level. For example, one work alters biased embeddings with minimal alterations to make them orthogonal to neutral embeddings (Rakshit et al., 2024). Additionally, some studies remove biases at the level of attention heads. Ma et al. (2023) address this by pruning biased attention heads. Beyond modifying embeddings and pruning attention heads, another strategy targets a specific group of neurons known to propagate biases. It unlearns the biases by retraining weight vectors for these neurons (Yu et al., 2023). Besides, bias mitigation can also be approached from a data-centric perspective using the most biased training samples, then modifying them to fine-tune the model (Thakur et al., 2023).
5.4. Toxicity
@@
Toxicity is another form of harmful content that LLMs may produce. This issue arises from training documents containing elements of toxicity that can hardly be fully eliminated. Toxicity can be identified by interpreting LLM components like the feed-forward layers and attention heads. For instance, recent work reveals how toxicity is represented within LLMs by identifying multiple vectors and relevant dimensions promoting toxicity within the MLP layers (Lee et al., 2024). Furthermore, the exploration of geometric structures in per-layer representations offers another way to detect toxicity. (Balestriero et al., 2023) extract input features from MLPs to describe the domain of prompts and classifying toxic remarks.
@@
Aforementioned insights also shed light on mitigation strategies. Motivated by the finding that toxicity can be reduced by manipulating relevant vectors, Lee et al. (2024) utilizes paired toxic and non-toxic samples to fine-tune models so that non-toxic content is promoted. By examining the changes in the parameter matrices during the fine-tuning process, it substantiates that even minor adjustments to these critical vectors can reduce toxicity. Built on the observation that LLMs’ representations are updated by outputs from attention layers (Elhage et al., 2021), another work attempts to reduce toxicity by identifying the “toxicity direction” and then adjusting representations in the opposite direction (Leong et al., 2023).
5.5. Truthfulness